-4(3b+4) + (13b -6)=0
Answer by ReadingBoosters(3246)

Starting with
-4(3b+4) + (13b -6) = 0
Distribute the -4
-12b - 16 + 13b - 6 = 0
Combine like terms
b - 22 = 0
b - 22 + 22 = 0 + 22
b = 22
Proof
-4((3)(22)+4) + ((13)(22)-6)
-4(66+4) + (286-6)
-4(70) + 280
-280 + 280 = 0